### v1

# Ricci almost solitons: complete examples

Abstract.
We construct new examples of various solitons as warped products. There are classes of complete Ricci almost solitons and complete Ricci-Bourguignon solitons that can be explicitly described in terms of elementary functions.
Key words and phrases:
Ricci soliton, Ricci almost soliton, Ricci-Bourguignon soliton, Bakry-Émery tensor, Hessian, warped product
2010 Mathematics Subject Classification:
primary 53C25; secondary 53C44, 83C05
1. Introduction and results
A Ricci soliton is a self-similar solution of the Ricci flow, i.e., a Riemannian metric $g$ satisfying the equation
${\rm Ric}+\frac{1}{2}\mathcal{L}_{X}g+\lambda g=0$
for the Lie derivative in the direction of some vector field $X$ and a constant $\lambda$. Any Einstein space is trivially a Ricci soliton with $X=0$. If in addition $X$ is assumed to be a non-trivial conformal vector field, then it immediately follows that the Ricci soliton is Einstein. This case is often discussed in the literature, see [11], [6]. Besides the standard sphere, compact Ricci solitons seem to be rare [4, 5]. Sophisticated analytic and topological methods have been employed for obtaining non-compact Ricci solitons [22], [3], for the approach by doubly warped products see [10].
If $X$ is the gradient of a function $f$ then it is called a gradient Ricci soliton, characterized by the equation
(1) ${\rm Ric}+\nabla^{2}f+\lambda g=0\,,$
where $\nabla^{2}f$ is the Hessian. If $\lambda$ is a non-constant function then one talks about an Ricci almost soliton or gradient Ricci almost soliton, respectively. Einstein spaces that are Ricci almost solitons with a nontrivial $X$ have to admit a nontrivial conformal vector field. These spaces are classified in a finite list [13], also in pseudo-Riemannian geometry [14].
A family $t\mapsto g(t)$ of Riemannian metrics satisfying the evolution equation
$\frac{\partial g}{\partial t}=-2\left(\textrm{Ric}-\rho Sg\right)$
for a constant $\rho\in\mathbb{R}$ is said to evolve by the Ricci-Bourguignon flow. Here we will consider two particular cases. For $\rho=0$ this is the Ricci flow, for $\rho=1/n$ the right hand side equals the traceless Ricci tensor. This flow was introduced by Bourguignon [2, 3.24].
Then a gradient Ricci-Bourguignon soliton is a Riemannian metric $g$ together with a smooth function $f$ such that the following equation holds for some constant $\lambda:$
(2) $\textrm{Ric}+\nabla^{2}f-\rho Sg=\lambda g.$
Ricci-Bourguignon gradient solitons are also named gradient $\rho$-Einstein solitons, if $\lambda$ in Equation (2) is allowed to be a function, then the metric is also called gradient Ricci-Bourguignon almost soliton. For $\lambda=0$ the gradient Ricci-Bourguignon (almost) soliton is called steady.
In the present article we focus on complete and non-compact solutions of Equation (1) and Equation (2) on warped products. We find new Ricci almost solitons of this type in Section 3, and in Theorem 4.3 we classify for $n\geq 3$ and $\rho=1/n$ all complete and non-compact steady Ricci-Bourguignon solitons of this type, which then have a Ricci flat fibre. All solutions are explicit in terms of elementary functions. Ricci-Bourguignon solitons as generalized warped products were investigated in [12].
Main result.
All complete warped products with a $1$-dimensional basis satifying the equation $\textrm{Ric}+\nabla^{2}f-\frac{S}{n}g=0$ of a steady Ricci-Bourguignon soliton can be classified by a family of explicitly given functions $h$ and $f$ depending on finitely many parameters.
Throughout this paper ${\rm Ric}$ and $({\rm Ric})^{\circ}={\rm Ric}-\frac{S}{n}g$ will denote the Ricci tensor and the traceless Ricci tensor, respectively, $S$ is the scalar curvature.
2. Basic formulas
We consider an $n$-dimensional Riemannian manifold $(M,g)\cong\mathbb{R}\times(M_{*},g)$ with $n\geq 2$ and with a warped product metric
$ds^{2}=dt^{2}+h^{\prime 2}(t)ds_{*}^{2}$
and a function $f(t)$ defined on the basis. Unless stated otherwise, it will be assumed that $h$ and $f$ are smooth functions and $h^{\prime}$ is not constant. Then the gradients of $h$ and $f$ appear as $h^{\prime}\partial_{t}$ and $f^{\prime}\partial_{t}$, respectively, and the Hessian tensors are $\nabla^{2}h=h^{\prime\prime}g$ on the one hand and
$\nabla_{\partial_{t}}\nabla f=f^{\prime\prime}\partial_{t}$
and
$\nabla_{X}\nabla f=\nabla_{X}\left(f^{\prime}\partial_{t}\right)=f^{\prime}\nabla_{X}\partial_{t}=\frac{f^{\prime}}{h^{\prime}}\nabla_{X}\nabla h=\frac{f^{\prime}h^{\prime\prime}}{h^{\prime}}X$
for any $X$ perpencidular to $\partial_{t}$ on the other hand.
The Ricci tensors of $g$ and $g_{*}$ will be denoted by Ric and ${\rm Ric}_{*}$. Any equation of the type
(3) $a\textrm{Ric}+b\nabla^{2}f+cg=0$
with functions $a,b,c$ on $M$ leads to equations of the form
$a\textrm{Ric}(\partial_{t},\partial_{t})+bf^{\prime\prime}+c$ $=$ 00 0
$a\textrm{Ric}(X,X)+b\frac{f^{\prime}h^{\prime\prime}}{h^{\prime}}g(X,X)+cg(X,X)$ $=$ $0\,.$
For the warped product metric above we obtain in direction $\partial_{t}:$
(4) $-(n-1)a\frac{h^{\prime\prime\prime}}{h^{\prime}}+bf^{\prime\prime}+c=0$
and in any direction $X$ perpendicular to $\partial_{t}:$
(5) $a\textrm{Ric}_{*}+\left\{-a(n-2)h^{\prime\prime 2}-ah^{\prime}h^{\prime\prime\prime}+bf^{\prime}h^{\prime}h^{\prime\prime}+ch^{\prime 2}\right\}g_{*}=0.$
From Equation (5) we see that $(M_{*},g_{*})$ is Einstein, so for $n-1\geq 3$ there is a constant $\lambda_{*}$ (the normalized Einstein constant of $M_{*}$) satisfying
(6) $a(n-2)h^{\prime\prime 2}+ah^{\prime}h^{\prime\prime\prime}-bf^{\prime}h^{\prime}h^{\prime\prime}-ch^{\prime 2}=\lambda_{*}a.$
For given constants $a,b$ and given $h^{\prime}$ one can easily determine an appropriate $f^{\prime}$ and a function $c$ such that the equations above are satisfied.
We eliminate $c$ from the equations by inserting Equation (4) into Equation (6):
(7) $a(n-2)h^{\prime\prime 2}+ah^{\prime}h^{\prime\prime\prime}-bf^{\prime}h^{\prime}h^{\prime\prime}-(n-1)ah^{\prime}h^{\prime\prime\prime}+bf^{\prime\prime}h^{\prime 2}=\lambda_{*}a$
or, equivalently,
(8) $a(n-2)\big{[}h^{\prime\prime 2}-h^{\prime}h^{\prime\prime\prime}\big{]}+bh^{\prime}\big{[}f^{\prime\prime}h^{\prime}-f^{\prime}h^{\prime\prime}\big{]}=\lambda_{*}a$
or
(9) $a(n-2)h^{\prime\prime 2}\Big{(}\frac{h^{\prime}}{h^{\prime\prime}}\Big{)}^{\prime}+bh^{\prime 3}\Big{(}\frac{f^{\prime}}{h^{\prime}}\Big{)}^{\prime}=\lambda_{*}a.$
Therefore one can express $f^{\prime}$ as $h^{\prime}$ times an integral of $\frac{a}{b}(n-2)\frac{h^{\prime\prime 2}-h^{\prime}h^{\prime\prime\prime}-\lambda_{*}}{h^{\prime 3}}$ and determine $f$ itself by an additional integration step, compare [18, Ex.1.5]. Finally, $c$ is obtained by the trace of $a{\rm Ric}+b\nabla^{2}f$. However, in general $c$ won’t be constant, so this method does not produce Ricci solitons but only Ricci almost solitons. Considering only the traceless part leads to the following:
Proposition 2.1.
For any given Einstein space $M_{*}$ and any smooth function $h^{\prime}$ there is a function $f$ such that the warped product ×$\mathbb{R}\times_{h^{\prime}}M_{*}$ admits at least locally a solution of $(a{\rm Ric}+b\nabla^{2}f)^{\circ}=0$. For $a=b=1$ the left hand side is nothing but the traceless Bakry-Émery-Ricci tensor [21].
If $h^{\prime}$ is globally defined and everywhere strictly positive, then the warped product ×$\mathbb{R}\times_{h^{\prime}}M_{*}$ is complete whenever $M_{*}$ is complete. This leads to complete solutions of $({\rm Ric}+\nabla^{2}f)^{\circ}=0$ (Ricci almost solitons).
Proof. Equation (8) expresses the fact that both Equation (4) and Equation (5) are simultaneously satisfied with a certain function $c(t)$. This implies Equation (3) for this unknown $c$ or, equivalently, $(a{\rm Ric}+b\nabla^{2}f)^{\circ}=0$. The second claim is obvious from the explicit formula for $f^{\prime}$ in terms of $h^{\prime}>0$. $\Box$
Proposition 2.2.
(special cases)
  1. (1)
    The case $a=1$ and $b=0$ and, consequently, $(n-2)\big{[}h^{\prime\prime 2}-h^{\prime}h^{\prime\prime\prime}\big{]}=\lambda_{*}$ leads to an Einstein space and the standard (non-constant) solutions $h^{\prime}(t)=\sin t,\cos t,\sinh t,\cosh t,e^{t},t$ of $(\nabla^{2}h)^{\circ}=0$ on such [13], [14].
  2. (2)
    Essentially the same holds under the assumption that $a=b=1$ and $M$ is Einstein, compare [18, Thm.1.3]. Up to scaling, the list of Einstein spaces admitting a function $f$ satisfying $\nabla^{2}f=\lambda g$ for some function $\lambda$ consists of warped products $dt^{2}+(f^{\prime}(t))^{2}ds_{*}^{2}$ with a $1$-dimensional basis where $f(t)$ is a solution of the oscillator equation $f^{\prime\prime}+\sigma f=0$ where $\sigma$ is a normalized scalar curvature ($\sigma=1$ for the unit sphere) [13], [14]. Among them the unit sphere is the only compact Einstein space with $f^{\prime}(t)=\sin t$ with the two zeros added (it is not a global warped product).
  3. (3)
    The case $a=0$ and $b=1$ leads to $f^{\prime}=\gamma h^{\prime}$ with a constant $\gamma$, the classical solutions of $(\nabla^{2}f)^{\circ}=0$, a conformal gradient field [13].
  4. (4)
    Conversally, the assumption $f^{\prime}=\gamma h^{\prime}$ with a constant $\gamma$ leads to $(\nabla^{2}f)^{\circ}=0$ and thus to an Einstein space if $a\not=0$.
Remark.
A solution with $a=b=1$ and a constant $c$ is normally called a gradient Ricci soliton, and the tensor ${\rm Ric}^{\infty}_{f}={\rm Ric}+\nabla^{2}f$ is called the Bakry-Émery-Ricci Tensor [21], for ${\rm Ric}^{m}_{f}$ see [9]. A solution with $a=b=1$ and a function $c$ is called a gradient Ricci almost soliton [18].
Proposition 2.3.
(The case of low dimensions)
  1. (1)
    Any $2$-dimensional gradient Ricci soliton with $b\not=0$ is a warped product outside the set of critical points which are isolated. This follows from the known solutions of the equation $\nabla^{2}f=\phi g$ with any scalar function $\phi$ [19], [13], compare [17].
  2. (2)
    Conversally, a $2$-dimensional warped product metric as above is a gradient Ricci soliton if it admits a function $f(t)$ such that $\nabla^{2}f=-(K+c)g$ where $K$ denotes the Gaussian curvature $K=-h^{\prime\prime\prime}/h^{\prime}$. This implies $f^{\prime\prime}=f^{\prime}h^{\prime\prime}/h^{\prime}=h^{\prime\prime\prime}/h^{\prime}-c$ and, consequently, $f^{\prime}=\gamma h^{\prime}$ with a constant $\gamma$. Locally this ODE has solutions for any $c$.
  3. (3)
    For $n=3$ the warped product approach with the equations above requires the extra assumption that $M_{*}$ is a space of constant Gaussian curvature $\lambda_{*}$.
  4. (4)
    For $n=4$ the warped product approach with the equations above implies that $M_{*}$ is a space of constant sectional curvature $\lambda_{*}$.
Examples 2.4.
1. The metric $ds^{2}=dt^{2}+\tanh^{2}t\,du^{2}$ on $(t,u)\in[0,\infty)\times S^{1}$ is a well known example of a $2$-dimensional complete Ricci soliton defining a complete metric on $\mathbb{R}^{2},$ cf. Kühnel [13, Example 4(a)]. Here $c=0,\gamma=-2,h^{\prime}(t)=\tanh t,f^{\prime}=-2h^{\prime},h^{\prime\prime}(t)=(\cosh t)^{-2},K=-h^{\prime\prime\prime}/h^{\prime}=2h^{\prime\prime}=-f^{\prime\prime},\nabla^{2}f=f^{\prime\prime}g=-2h^{\prime\prime}g=K\,g.$ With $x=\sinh t\,\sin u;y=\sinh t\cos u$ this metric on $\mathbb{R}^{2}$ is of the form
$ds^{2}=\frac{dx^{2}+dy^{2}}{1+x^{2}+y^{2}};f(x,y)=-\log(1+x^{2}+y^{2}),$
it is called Hamilton’s cigar soliton in the literature, cf. [8].
2. The product of an Euclidean $\mathbb{R}^{m}$ with a $k$-dimensional Einstein space with Einstein constant $\lambda\not=0$ is a gradient Ricci soliton for the function $f(x)=\lambda|x|^{2}$ and for $c=-\lambda$.
3. Complete and explicit solutions
Complete solutions are of particular interest. Proposition 2.1 gives a possibility for constructing Ricci almost solitons as warped products. In fact there are complete Ricci almost solitons that can be explicitly described in terms of elementary functions. For $n=2$ the situation is better because one can prescribe the three constants $a,b,c$ and solve the resulting ODE for one function (not coupled as in the case $n\geq 3$).
Proposition 3.1.
If $n=2$ then the two coupled equations (4) und (6) read as:
$-ah^{\prime\prime\prime}+bh^{\prime}f^{\prime\prime}+ch^{\prime}$ $=$ 00 0
$ah^{\prime}h^{\prime\prime\prime}-bf^{\prime}h^{\prime}h^{\prime\prime}-ch^{\prime 2}$ $=$ 00 0
For $h^{\prime}\not=0$ these are equivalent to the following two uncoupled equations:
(10) $f^{\prime}$ $=$ $\gamma h^{\prime}$
(11) $-ah^{\prime\prime\prime}+b\gamma h^{\prime}h^{\prime\prime}+ch^{\prime}$ $=$ 00 0
for some constant $\gamma\not=0$ (the case $\gamma=0$ is trivial). The second equation allows explicit solutions at least for the inverse function $t=t(h^{\prime})$.
Proof. Both equations are simultaneously satisfied if and only if either $b=0$ or $h^{\prime}f^{\prime\prime}=h^{\prime\prime}f^{\prime}$ which in turn is equivalent to $\log(f^{\prime})^{\prime}=\log(h^{\prime})^{\prime}$ or $f^{\prime}=\gamma h^{\prime}$ for a constant $\gamma$. $\Box$
There is a method for solving an ODE of type $y^{\prime\prime}=F(y,y^{\prime})$ [20, §11]. We concentrate on the case $a=b=1$.
Lemma 3.2.
For $c=0$ Equation (11) is of first order by $(h^{\prime\prime}-\frac{\gamma}{2}h^{\prime 2})^{\prime}=0$. There are various odd functions $h^{\prime}$ solving this equation. Typical solutions with initial condition $h^{\prime}(0)=0$ are $h^{\prime}(t)=\tan t$ for $\gamma=2$ and $h^{\prime}(t)=\tanh t=\frac{e^{2t}-1}{e^{2t}+1}$ for $\gamma=-2$ (the complete cigar metric. Other typical solutions are $h^{\prime}(t)=\frac{1}{t}$ for $\gamma=-2$ and $h^{\prime}(t)=\coth t=\frac{e^{2t}+1}{e^{2t}-1}$ for $\gamma=-2$. There are no even functions solving this equation.
Proof. In the first case we have $h^{\prime\prime}-h^{\prime 2}=1$, in the second case $h^{\prime\prime}+h^{\prime 2}=1$, in the third case $h^{\prime\prime}+h^{\prime 2}=0$, in the fourth case $h^{\prime\prime}+h^{\prime 2}=1$.
Example 3.3.
Let $h(t)=\arctan(t)$, then for any $\lambda_{*}$ there is a global function $f(t)$ such that ×$\mathbb{R}\times_{h^{\prime}}M_{*}$ is a complete Ricci almost soliton whenever $M_{*}$ is complete with Einstein constant $\lambda_{*}$.
Calculation. We have $h^{\prime}(t)=\frac{1}{t^{2}+1}>0$ and $h^{\prime\prime}(t)=-\frac{2t}{(t^{2}+1)^{2}}$. For any constants $a,b\not=0$ we have to solve Equation (9):
$a(n-2)\big{[}h^{\prime\prime 2}-h^{\prime}h^{\prime\prime\prime}\big{]}+bh^{\prime 3}\Big{(}\frac{f^{\prime}}{h^{\prime}}\Big{)}^{\prime}=-a(n-2)\frac{2(t^{2}-1)}{(t^{2}+1)^{4}}+b\frac{1}{(t^{2}+1)^{3}}\big{(}(t^{2}+1)f^{\prime}\big{)}^{\prime}=a\lambda_{*}.$
This leads to
$\big{(}(t^{2}+1)f^{\prime}\big{)}^{\prime}=\frac{a\lambda_{*}}{b}(t^{2}+1)^{3}+\frac{a(n-2)}{b}\cdot\frac{2(t^{2}-1)}{t^{2}+1}$
and
[…]
$=\frac{a\lambda_{*}}{b}\Big{(}\frac{t^{7}}{7}+\frac{3t^{5}}{5}+t^{3}+t\Big{)}+\frac{a(n-2)}{b}\Big{(}2t-4\arctan(t)+C\Big{)}.$
For obtaining $f(t)$ itself it only remains to evaluate the integral in
$f(t)$ $=$ $\frac{a\lambda_{*}}{b}\int\Big{(}\frac{t^{7}}{7}+\frac{3t^{5}}{5}+t^{3}+t\Big{)}\frac{dt}{t^{2}+1}$
$+\frac{a(n-2)}{b}\Big{(}\log(t^{2}+1)-2\big{(}\arctan(t)\big{)}^{2}+C\arctan(t)+D\Big{)}$
with constants $C,D$. All expressions are globally defined on $\mathbb{R}$ and give a global solution with $h^{\prime}(t)\not=0$.
[…]
$a({\rm Ric})^{\circ}+b(\nabla^{2}f)^{\circ}=0\ \mbox{ or }\ a{\rm Ric}+b\nabla^{2}f+cg=0$
with $c=-(aS+b\Delta f)/n$.
Example 3.4.
Let $h(t)=\frac{1}{3}t^{3}+t$, then for any $\lambda_{*}$ there is a global function $f(t)$ such that ×$\mathbb{R}\times_{h^{\prime}}M_{*}$ is a complete Ricci almost soliton whenever $M_{*}$ is complete with Einstein constant $\lambda_{*}$.
Calculation. We have $h^{\prime}(t)=t^{2}+1>0$ and $h^{\prime\prime}(t)=2t$. This implies $h^{\prime\prime 2}-h^{\prime}h^{\prime\prime\prime}=4t^{2}-2(t^{2}+1)=2(t^{2}-1)$. Therefore we have to solve the equation
$\Big{(}\frac{f^{\prime}}{t^{2}+1}\Big{)}^{\prime}=\Big{(}\lambda_{*}a-2a(n-2)(t^{2}-1)\Big{)}\frac{1}{b(t^{2}+1)^{3}},$
and $f^{\prime}(t)$ can be determined by integration of rational functions. In a second step one can express $f(t)$ at least by integrals of rational functions and their integrals. Obviously $f(t)$ is defined for any $t\in\mathbb{R}$
If $g_{*}$ is the standard metric on the sphere, then the space can be viewed as a rotationally symmetric deformed hyperboloid which opens up quadratically at the two ends.
Example 3.5.
Let $h(t)=(t^{2}+2)e^{t}$, then for any $\lambda_{*}$ there is a global function $f(t)$ such that ×$\mathbb{R}\times_{h^{\prime}}M_{*}$ is a complete Ricci almost soliton whenever $M_{*}$ is complete with Einstein constant $\lambda_{*}$.
Calculation. We have $h^{\prime}(t)=(t^{2}+2t+2)e^{t}>0,h^{\prime\prime}(t)=(t+2)^{2}e^{t},h^{\prime\prime\prime}(t)=(t^{2}+6t+8)e^{t}$. We have to solve the equation
$a(n-2)\big{[}h^{\prime\prime 2}-h^{\prime}h^{\prime\prime\prime}\big{]}+bh^{\prime 3}\Big{(}\frac{f^{\prime}}{h^{\prime}}\Big{)}^{\prime}$
$=-2a(n-2)(t^{2}+12t+8)e^{2t}+b(t^{2}+2t+2)^{3}e^{3t}\Big{(}\frac{f^{\prime}}{(t^{2}+2t+2)e^{t}}\Big{)}^{\prime}=a\lambda_{*}.$
$f^{\prime}$ can be determined by explicit integration, $f$ can at least be expressed by integrals.
If $g_{*}$ is the standard metric on the sphere, then the space can be viewed as a rotationally symmetric deformed infinite trumpet with a cusp at one end and a huge opening at the other.
Remark (The power series approach for Ricci solitons).
For $a=b=1$ and a constant $c$ one can try to solve the system of equations 4 und 6 simultaneously by a power series approach:
$h^{\prime}(t)=\sum_{i\geq 0}a_{i}t^{i},\quad f^{\prime}(t)=\sum_{j\geq 0}b_{j}t^{j}$
For $c=-1,a_{0}=b_{0}=0,a_{1}=b_{1}=1$ and $\lambda_{*}=n-2$ we have the classical solution (Gaussian Ricci soliton) with $h^{\prime}(t)=f^{\prime}(t)=t$. This is nothing but the Euclidean space in polar coordinates with the function $f(x)=\frac{1}{2}|x|^{2}$, compare Proposition 2.2(2). Under the assumption that both $h^{\prime}$ and $f^{\prime}$ are odd functions, one can transform Equation (4) into a (very complicated) recursion formula for the coefficients depending on given constants $a_{1},b_{1},c$ and $\lambda_{*}$. Note, however, that even for the complete cigar solution in Examples 2.4 (1) the power series has a finite radius of convergence.
4. Ricci-Bourguignon solitons
An $n$-dimensional Ricci-Bourguignon soliton is a self-similar solution of the flow with any tensor ${\rm Ric}-\rho Sg$ instead of the Ricci tensor, where $\rho$ is a constant [2, 3.24], [6]. The case $\rho=\frac{1}{2(n-1)}$ is referred to also as a Ricci-Schouten soliton [1]. In particular we are going to deal with the traceless Ricci tensor instead of the Ricci tensor itself, so we have $\rho=\frac{1}{n}$. We are going to investigate solutions of the equation
(12) ${\rm Ric}+\nabla^{2}f-\frac{S}{n}g=\lambda g$
with a constant $\lambda$ where $S$ denotes the scalar curvature [6]. The case $\lambda=0$ is then referred to as a steady Ricci-Bourguignon gradient soliton.
Remarks.
1. The only complete Einstein spaces satisfying Equation (12) are the Euclidean space in polar coordinates (if $\nabla^{2}f=\lambda g$ with a constant $\lambda\not\equiv 0$) and products of an interval with a complete Ricci flat manifold (if $\nabla^{2}f=0$). This follows from the classification of Einstein spaces admitting a solution of the equation $\nabla^{2}f=\phi g$ with a scalar function $\phi$ [13], [14].
2. If the gradient of $f$ is assumed to be a conformal vector field, then Equation (12) implies that $M$ is Einstein. Consequently, even if $\lambda$ is admitted to be a non-constant function, the only complete cases are listed in [13, Thm.27], implicitly already in [19, Sect.3]. In particular, the only compact case is the standard sphere. Implicitly this result can already be found in [19, Thm.1] since an Einstein space conformal to the standard sphere is isometric to the standard sphere. It is restated in [6, Thm.1.5]. For the last statement a constant scalar curvature is sufficient [13, Thm.24].
3. Complete gradient Ricci-Bourguignon almost solitons can be constructed by the same principle as in Section 3, compare [6, Ex.1.7].
4. The case $n=2$ is not interesting here since the traceless Ricci tensor always vanishes.
Proposition 4.1.
A warped product $M=\mathbb{R}\times_{h^{\prime}}M_{*}$ of dimension $n\geq 4$ with a potential function $f(t)$ is a steady Ricci-Bourguignon gradient soliton if and only if $f$ is harmonic and Equation (8) is satisfied with $a=b=1$.
The same holds for $n=3$ under the assumption that $g_{*}$ is a metric of constant curvature $\lambda_{*}$.
Proof. From the trace of Equation (12) we obtain $\Delta f=\lambda=0$. Furthermore Equation (8) holds in general for this type of warped products, as shown in Section 2.
Conversely, Equation (8) with $a=b=1$ implies ${\rm Ric}+\nabla^{2}f+cg=0$ for some function $c$. From the trace of this equation and $\Delta f=0$ one sees that $c=-\frac{S}{n}$. $\Box$
Corollary 4.2.
An $n$-dimensional warped product metric
$ds^{2}=dt^{2}+(h^{\prime}(t))^{2}ds_{*}^{2}$ is a steady Ricci-Bourguignon gradient soliton if and only if $h$ and $f$ satisfy the following equations with the constant $\lambda_{*}$ as the normalized scalar curvature of $g_{*}$:
(13) $(n-2)\big{[}h^{\prime\prime 2}-h^{\prime}h^{\prime\prime\prime}\big{]}+h^{\prime}\big{[}f^{\prime\prime}h^{\prime}-f^{\prime}h^{\prime\prime}\big{]}$ $=$ $\lambda_{*}$
(14) $\Delta f=f^{\prime\prime}+(n-1)\frac{f^{\prime}h^{\prime\prime}}{h^{\prime}}$ $=$ $0$
Theorem 4.3.
All complete warped products in dimension $n\geq 3$ according to Proposition 4.1 and with $\lambda_{*}=0$ can be classified as follows:
$(M_{*},g_{*})$ must be complete and Ricci flat, the possible non-constant functions $h^{\prime}>0$ and $f^{\prime}$ are the following:
$h^{\prime}(t)=\Big{(}\big{[}\frac{B}{D}+E^{n-1}\big{]}e^{(n-1)Dt}-\frac{B}{D}\Big{)}^{1/(n-1)}\ \ \mbox{ and }\ \ f^{\prime}=C(h^{\prime})^{1-n}$
with constants $B>0,C>0,D<0,E=h^{\prime}(0)$ such that $E^{n-1}>-\frac{B}{D}$.
Proof. The case of an Einstein space can be ruled out here since there are only the examples with a constant $h^{\prime}$, compare the Remark 1 above. The Euclidean space in polar coordinates with $h^{\prime}=t$ and $\lambda_{*}=1$ is not an example because $\nabla^{2}f=0$ is not possible in this case for a function $f(t)$ although there is a harmonic function $f(t)$. Now we assume that $h^{\prime}$ is not constant and is nowhere zero. From Corollary 4.2 we obtain that the two equations
$(n-2)\big{[}h^{\prime\prime 2}-h^{\prime}h^{\prime\prime\prime}\big{]}+h^{\prime}\big{[}f^{\prime\prime}h^{\prime}-f^{\prime}h^{\prime\prime}\big{]}$ $=$ 00 0
$f^{\prime\prime}+(n-1)\frac{f^{\prime}h^{\prime\prime}}{h^{\prime}}$ $=$ $0$
are satisfied. The second equation is equivalent to
$\frac{f^{\prime\prime}}{f^{\prime}}=-(n-1)\frac{h^{\prime\prime}}{h^{\prime}}\ \mbox{ or }\ \log(f^{\prime})=-(n-1)\log(h^{\prime})+C_{0}$
with a constant $C_{0}$, so it is equivalent to
(15) $f^{\prime}=C(h^{\prime})^{1-n}$
with a constant $C>0$. Therefore $f$ is globally defined whenever $h^{\prime}$ is globally defined without a zero. In a zero of $h^{\prime}$ the function $f^{\prime}$ tends to infinity, so it is really a singularity. This case will occur, depending on the choice of an additional constant $D$ that comes in.
If we insert the second into the first equation we obtain
$(n-2)\big{[}h^{\prime\prime 2}-h^{\prime}h^{\prime\prime\prime}\big{]}+h^{\prime}\big{[}-nf^{\prime}h^{\prime\prime}\big{]}=0$
or, equivalently
(16) $(2-n)\Big{(}\frac{h^{\prime\prime}}{h^{\prime}}\Big{)}^{\prime}+\big{[}-nCh^{\prime\prime}(h^{\prime})^{-n}\big{]}=0.$
We can solve Equation (16) by
$h^{\prime\prime}(h^{\prime})^{-n}=\frac{1}{1-n}\Big{(}(h^{\prime})^{1-n}\Big{)}^{\prime}$
which implies
$(2-n)\frac{h^{\prime\prime}}{h^{\prime}}=\frac{nC}{1-n}(h^{\prime})^{1-n}+D(2-n)$
with another constant $D$. Now the discussion splits into the two cases
$h^{\prime\prime}=B(h^{\prime})^{2-n}$
with the constant $B=\frac{nC}{(n-1)(n-2)}>0$  if $D=0$  and
$h^{\prime\prime}=B(h^{\prime})^{2-n}+Dh^{\prime}$
otherwise.
1. The case $D=0$
Here the solution is
$h^{\prime}(t)=\Big{(}B(n-1)t+E^{n-1}\Big{)}^{\frac{1}{n-1}}$
with an additional constant $E$ according to the initial condition $h^{\prime}(0)=E$. Necessarily $h^{\prime}$ has a zero at $t=t_{0}$ for a certain $t_{0}$. Therefore this solution does not lead to a complete manifold. The singularity at $t=t_{0}$ is recognized by an infinite value of $h^{\prime\prime}$ and also of $f^{\prime}=C\big{(}h^{\prime}\big{)}^{1-n}$. The qualitative behavior does not depend on the choice of the constant $E$.
2. The case $D\not=0$
The general solution is
$h^{\prime}(t)=\Big{(}\big{[}\frac{B}{D}+E^{n-1}\big{]}e^{(n-1)Dt}-\frac{B}{D}\Big{)}^{1/(n-1)}$
with an additional constant $E$ according to the initial condition $h^{\prime}(0)=E$. Obviously this solution $h^{\prime}(t)$ is globally defined and strictly positive if and only if $D<0$ and $E^{n-1}>-\frac{B}{D}>0$. Recall $B>0$ in any case. In the limit $t\to\infty$ the metric is asymptotic to the product metric $dt^{2}+\big{|}\frac{B}{D}\big{|}^{2/(n-1)}ds_{*}^{2}$ on ×$\mathbb{R}\times M_{*}$ which is Ricci flat. Strictly negative solutions can be considered only if $n$ is even but these cases don’t give any really new solutions for the warped product metric. $\Box$
Remarks. 1. For fixed $B$ and $E$ the transition $D\to 0$ in the second case leads to the limits $\lim_{D\to 0}B\frac{e^{(n-1)Dt}-1}{D}=B(n-1)t$ and $\lim_{D\to 0}e^{(n-1)Dt}=1$ for any $t$ (pointwise), so the limit function coincides with the solution in the case $D=0$. The singularity (a zero of $h^{\prime}$) occurs if $E^{n-1}<-\frac{B}{D}$, and $h^{\prime}$ is constant for $E^{n-1}=-\frac{B}{D}$.
2. The global shape in the case $D\not=0$ can look like a huge trumpet with a compact Ricci flat space as the fibre. with a cusp at one end and an infinite opening at the other. In the case $D=0$ it looks like a cigar, modelled by the function $h^{\prime}(t)=\sqrt[n-1]{t}$. If $M_{*}$ is compact and Ricci-flat then there is a topological singularity at $t=0$.